3a over 4a-20 + 9a over 6a-30
show work

Answer by fractalier(6550)

To combine these, we must convert each to their lowest common denominator...in this case that would be 12a-60...so we have
3a / (4a-20) + 9a / (6a-30) =
9a / (12a-60) + 18a / (12a-60) =
27a / (12a-60) now reduce to get
27a / 12(a-5) =
9a / 4(a-5)